OpenStudy (anonymous):

log6x+log6 (x-1) = 1

OpenStudy (australopithecus):

log_6(x(x-1)) = 1 6^(log_6(x(x-1))) = 6^(1) x(x-1) = 6

OpenStudy (anonymous):

Distribute the log6 to (x-1)

OpenStudy (anonymous):

combine like terms and then isolate x

OpenStudy (australopithecus):

Romero I don't think your advice is useful, no offense, combine logs and then elminate the logirithym by turning both sides of the equations to exponents of 6

OpenStudy (australopithecus):

List of log rules here if you dont understand how I simplified this http://openstudy.com/users/australopithecus#/updates/4f8a2358e4b02251ecca446b